On January 1, 2021, Company A leased equipment from Company B. The lease agreement specifies seven annual payments of $50,000 beginning January 1, 2021, the beginning of the lease, and at each December 31, thereafter. At the end of the seven-year lease term the equipment will be returned to the lessor and is expected to have a residual value of $24,000. The estimated useful life of the equipment is eight years. The interest rate in these financing arrangements is 10%. Company B manufactured the equipment at a cost of $200,000. Which of the following is true regarding the entries made on January 1, 2021?

Group of answer choices

Company B will debit equipment for $200,000.

Company B will credit sales revenue for $267,763.

Company B will credit sales revenue for $280,079.

Company B will debit lease receivable for $267,763.
Answer #1

Answer: Company B will credit sales revenue for $267,763.

Calculation

Year PV factor @ 8% Remarks
0    1.00000
1    0.90909 = 1 / 1.10
2    0.82645 = 0.90909 / 1.10
3    0.75131 = 0.82645 / 1.10
4    0.68301 = 0.75131 / 1.10
5    0.62092 = 0.68301 / 1.10
6    0.56447 = 0.62092 / 1.10
7    0.51316 = 0.56447 / 1.10
Total (0 to 6)    5.35526
For Lessor (Company B)
Present value of lease payments (50000*5.35526) $        267,763
Present value of residual value (24000*0.51316) $          12,316
Lease receivable $        280,079
Present value of unguaranteed residual value ($12,316) is deducted from cost of goods sold and sales revenue.
Date Accounts titles and explanation Debit Credit
January 1, 2021 Lease receivable    280,079
Cost of goods sold (200000-12316)    187,684
Sales revenue (280079-12316)            267,763
Equipment            200,000
(To record lease agreement for sales type lease.)